# Learning from Reasoning Failures via Synthetic Data Generation

### v1

@@ Abstract
Training models on synthetic data has emerged as an increasingly important strategy for improving the performance of generative AI. This approach is particularly helpful for large multimodal models (LMMs) due to the relative scarcity of high-quality paired image-text data compared to language-only data. While a variety of methods have been proposed for generating large multimodal datasets, they do not tailor the synthetic data to address specific deficiencies in the reasoning abilities of LMMs which will be trained with the generated dataset. In contrast, humans often learn in a more efficient manner by seeking out examples related to the types of reasoning where they have failed previously. Inspired by this observation, we propose a new approach for synthetic data generation which is grounded in the analysis of an existing LMM’s reasoning failures. Our methodology leverages frontier models to automatically analyze errors produced by a weaker LMM and propose new examples which can be used to correct the reasoning failure via additional training, which are then further filtered to ensure high quality. We generate a large multimodal instruction tuning dataset containing over 553k examples using our approach and conduct extensive experiments demonstrating its utility for improving the performance of LMMs on multiple downstream tasks. Our results show that models trained on our synthetic data can even exceed the performance of LMMs trained on an equivalent amount of additional real data, demonstrating the high value of generating synthetic data targeted to specific reasoning failure modes in LMMs. We will make our dataset and code publicly available.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have significantly advanced the state-of-art across a broad range of problem domains, which can be attributed in large part to scaling the size of models and datasets used for training. However, a limiting factor to further scaling of training datasets is the availability of new high-quality data. With recent studies suggesting that we are rapidly approaching saturation of existing public data sources [39, 42], synthetic data generation has become an increasingly popular approach for augmenting training datasets. The generation of high-quality synthetic data has become more feasible recently as the capabilities of LLMs have increased, thereby enabling their use for the production of synthetic data at scale [7, 45, 20, 13, 14, 35].
Synthetic data has been particularly valuable for training large multimodal models (LMMs), which combine an LLM with a vision encoder to enable text generation conditioned on multimodal inputs. In contrast to LLMs, real data suitable for training LMMs is relatively scarce due to the lack of naturally-occurring images paired with the type of high-quality text expected to be produced by such models. Consequently, popular LMMs such as LLaVA [22] use partially synthetic data for multimodal instruction tuning, which is often collected by generating synthetic text for real images.
@@
In contrast to the dominant paradigm for generating synthetic training data, humans often learn more efficiently by focusing their attention on examples which are most related to their past reasoning failures. To acquire expert knowledge, Ericsson et al. [8] argue that humans should “ideally be given explicit instructions about the best method and be supervised by a teacher to allow individualized diagnosis of errors, informative feedback, and remedial part training.” Since humans acquire new knowledge by learning from their failures [6, 3, 49, 30], seeking out additional examples related to past reasoning errors for further practice and evaluation helps the learner achieve expertise in a subject. Problems which require types of reasoning that have already been mastered are often ignored under this paradigm, as they provide little incremental value in the learning process.
@@
Using our approach, we generate a large multimodal instruction tuning dataset containing over 553k examples derived from the reasoning failures of LLaVA-1.5-7B, a popular LMM. We conduct extensive training experiments showing that our synthetic data improves LLaVA’s performance on multiple downstream tasks, even outperforming additional training on alternative real datasets. This is notable, as prior studies have shown that training on synthetic data often requires significantly more examples to achieve equivalent results as training on real data [12]. Training on our dataset also produces greater performance gains than previously-proposed synthetic datasets, highlighting the utility of grounding synthetic data generation in the analysis of a model’s reasoning failures.
@@
  5. 5.
    We make our dataset and code publicly available upon publication to support future research.
@@ 2 Related Work
Chen et al. [5] proposed the ALLaVA dataset, which consists of 1.3M samples of real images paired with image annotations and question-answer pairs generated by a frontier LMM. While their dataset is similarly suited for multimodal instruction tuning as our dataset, they utilize only real images and do not ground data generation in model failures, whereas our approach can utilize both real and synthetically generated images. Li et al. [21] generate synthetic question-answer pairs from an LLM for real chart images to address the narrower domain of reasoning-based chart VQA. Yang et al. [44] generate synthetic multimodal data for text-rich images, leveraging code-guided generation in languages such as LaTeX and HTML to automatically generate synthetic images with dense text descriptions. In contrast, our approach can be broadly applied to different problem domains while leveraging text-to-image diffusion models to produce more diverse types of images.
@@
Some prior work has explored the use of model failures in the context of synthetic data generation. Jain et al. [15] proposed an approach to identify directions in a vision model’s latent space which correspond to model failures, which can then be integrated into diffusion models to generate synthetic images tailored to correcting the model’s mistakes. Chegini and Feizi [4] utilize ChatGPT and CLIP to identify text descriptions corresponding to a vision model’s failure models, which are used as prompts for diffusion models to produce synthetic images for data augmentation. In the realm of language-only models, DISCERN [28] utilizes two LLMs to iteratively produce natural language descriptions of errors for synthetic data generation. Lee et al. [20] use incorrect answers from a student LLM finetuned on specific tasks as input to a teacher LLM which generates new examples to use for training. These prior works differ from ours in that they (1) generate only single-modality data (e.g., text-only or image-only datasets), and (2) focus on classification tasks such as image recognition or text classification. In contrast, our approach generates multimodal image-text datasets aimed at training models for open-ended text generation conditioned on multimodal inputs.
@@
AgentInstruct [29] is an agentic framework for generating synthetic data from a powerful frontier model (e.g., GPT-4) to teach new skills to a weaker LLM. Similarly, Ziegler et al. [48] utilize few-shot examples annotated by humans and retrieved documents with produce synthetic data from LLMs for teaching specialized tasks to models. Prompt-based methods for synthetic data generation from LLMs without seed documents [10, 37] as well as knowledge distillation from a teacher model [16] have also been proposed. Unlike our work, these prior studies focus on language-only data generation and use seed documents (e.g., raw text, source code) or prompts as a basis for data generation rather than an analysis of model failures.
@@ 3.2 Synthetic data generation
Using the resulting MFS, we guide the frontier model through a structured multi-turn process to diagnose the failure reasons of the baseline LMM and generate new training samples that address these failure modes, as illustrated in Figure 5. In the first steps, the frontier model is prompted to describe the image and analyze reasoning errors by examining the question, ground truth answer, and the incorrect response generated by the baseline LMM. Next, the frontier model is instructed to propose new challenging samples which consist of a detailed image description, a clear question, and a deterministic answer designed to target the identified failure modes. We explore two different approaches for sourcing images corresponding to these samples: one which utilizes existing real images, and another which uses synthetically generated images.
@@
We leverage the original images from failed samples and prompt the frontier model to generate 10 new question-answer pairs per sample using the reasoning pipeline described above. This approach is particularly effective for datasets that require precise text rendering or spatial accuracy, such as InfoVQA and ScienceQA. For these types of benchmarks, text-to-image models often struggle to generate coherent images that accurately reflect the textual content or capture the reasoning embedded in the prompt. This method utilizes the prompt described in Figure 5 (Section 7.3 of the supplementary material), but with the final step omitted.
@@
In this approach, the frontier model generates both a question-answer pair and a detailed image prompt, which is used to prompt a text-to-image diffusion model to generate synthetic images. For each failed sample, we create 100 fully synthetic samples, consisting of 10 (image prompt, question, answer) triplets, with each image prompt generating 10 images using different classifier-free guidance scales to enhance image diversity. This method utilizes the complete prompt depicted in Figure 5 (Section 7.3 of the supplementary material).
@@
To further enhance the diversity of our data, we use a variation of our prompt which instructs the frontier LMM to “provide examples that challenge Model A’s weaknesses using scenarios from entirely different domains or situations than the original example.” This additional instruction is appended to Step 4 (Figure 5) to encourage the generation of samples in a different domain. Figure 2 compares fully synthetic domain similar and non-similar samples. It is notable that domain-similar samples preserve the original theme, while non-similar samples cover a more diverse contextual range to improve generalization. Additionally, we created samples where we both enforced and relaxed constraints on question format (e.g., multiple-choice, true/false) and instructions (e.g., requiring responses like ”Unanswerable” when information was insufficient or limiting answers to short responses, see the Shiba Inu example from Figure 9).
@@
3.3 Filtering
We apply a filtering process using the same frontier LMM which produced the samples. The LMM is instructed to evaluate each synthetic sample given an image, question, and answer on a scale of 1 to 3, where 1 indicates an incorrect sample, 2 being partially correct, and 3 fully correct. The prompt used for filtering is provided in Figure 6 of Section  7.4. Only samples rated 3 were included in the final dataset to ensure quality and reliability.
@@ 3.4 Details of generated dataset
Our synthetic dataset consists of 553,992 samples incorporating both real and generated images derived from the MFS of LLaVA-1.5-7B on four benchmark training sets: VizWiz [11], InfoVQA [27], ScienceQA [25], and OK-VQA [26]. These benchmarks were chosen to ensure a wide range of visual and reasoning challenges. VizWiz consists of real-world images captured by visually impaired users, often requiring detailed scene understanding. OK-VQA focuses on visual question answering which requires external knowledge. InfoVQA involves text-rich images where reading comprehension is crucial, assessing the model’s ability to extract and interpret textual information from images. ScienceQA includes multimodal scientific reasoning questions which require both spatial and logical reasoning, making it valuable for evaluating complex reasoning capabilities. We use GPT-4o [31] as the frontier LMM for analyzing the reasoning failures of LLaVA-1.5-7B due to its strong multimodal reasoning capabilities. To generate the synthetic images, we utilized FLUX.1-schnell Labs [18] text-to-image model with a resolution of 1024×1024 pixels and a classifier guidance scale ranging from 3 to 13.
Dataset LLAVA $\textrm{LLaVA}_{\textrm{syn}}$
CIFAR-10  [17] 82.1 81.2
Food-101  [2] 13.4 13.2
iNaturalist  [38] 20.6 52.0
MNIST  [19] 75.1 80.5
F-MNIST  [41] 9.8 10.0
Oxford-pets  [32] 39.6 96.4
Table 1: Image classification accuracy of LLaVA and a $\textrm{LLaVA}_{\textrm{syn}}$ model augmented only with synthetic examples corresponding to object recognition failures.
@@
Table 5 (Section 7.2 of the supplementary material) provides an overview of each benchmark, including the original size of the training split, the MFS size of LLaVA-1.5-7b, and the total number of synthetic samples retained after filtering. As described in Section 3.2, we utilized two methods for synthetic data generation, depending on the dataset: for InfoVQA and ScienceQA, we applied Method 1, which generates new question-answer pairs based on the original real images. For OK-VQA, we employed Method 2, generating synthetic images along with corresponding question-answer pairs. For VizWiz dataset we integrated both methods, generating additional synthetic images as well as new question-answer pairs based on its existing real images. Overall, our final dataset consists of 42% real-image-based samples and 58% fully synthetic samples. All samples in the final dataset contain a single-turn conversation, featuring a question and a short answer.
Our filtering approach effectively removes lower-quality synthetic samples. For VizWiz, approximately 81% of synthetic-image samples and 34% of real-image-based samples were removed. This indicates, as expected, that generating entirely synthetic samples is more challenging than generating synthetic text alone for real images. OK-VQA had a lower removal rate of 29% for synthetic images compared to VizWiz, suggesting better synthetic image generation quality, possibly resulting from simpler and less ambiguous visual content. Among real-image-based samples, ScienceQA experienced a similar removal rate (29%), likely due to the complexity of spatial and scientific reasoning tasks. In contrast, InfoVQA exhibited a significantly lower removal rate of only 5% with an average filtering score of 2.9 (out of 3), indicating the strong capability of GPT-4o in handling text-based images. Nevertheless, the large number of generated data that passed filtering ensures comprehensive coverage of the reasoning failures previously identified in LLaVA-1.5-7b across various benchmarks. Figures 3 show sampled examples from VizWiz and InfoVQA highlighting the diversity of question types and demonstrating the overall quality of generated images and text. Appendix 10 provides additional examples from our dataset, including instances that are incorrect or ambiguous.
@@ 4.1 Details of experimental setting
We perform vision-language instruction tuning with LLaVA-v1.5-7B, building upon the LLaVA-1.5-mix-665K instruction tuning dataset [23]. This dataset contains 665K structured user-GPT conversations mostly focused on visual prompts. Because our approach explicitly targets image-grounded reasoning, we only selected the 624K samples that include images from the original LLaVA-instruct data. We then augmented this visually oriented subset with our generated synthetic MFS (Section 3.4), varying the number of synthetic samples used according to the requirements of each experiment. The resulting training set combines real image-text conversations and generated synthetic samples that address specific reasoning failures.
For fine-tuning, we used Vicuna-v1.5-7B [47] weights as the LLM backbone and leveraged the pretrained multimodal projector from LLaVA-1.5-7b. We followed the original LLaVA training procedure, ensuring a fair comparison to existing methods (see hyperparameters definition in Appendix 7.5). We refer to models trained with a mixture of the original LLaVA-Instruct dataset and our synthetic data as $\textrm{LLaVA}_{\textrm{syn}}$. As a baseline, we report the performance of LLaVA trained under the same setting but without any additional synthetic data added to the training dataset (i.e., $N_{syn}=0$). Additionally, for each dataset from which reasoning failures were derived for synthetic data generation, we report the performance of a LLaVA model trained on an equivalent amount of real data sourced from the corresponding training dataset (denoted as $\textrm{LLaVA}_{\textrm{real}}$). This provides a measure of the efficiency of our synthetic data relative to training on real in-domain data.
We used a variety of multimodal reasoning benchmarks for evaluating models. For in-domain evaluations (i.e., evaluating on a withheld validation set corresponding to the training set from which reasoning failures were derived), we used the InfoVQA, OK-VQA and ScienceQA validation sets. Since InfoVQA and ScienceQA questions rely heavily on reading text contained in the images, we further evaluated models on TextVQA [34] and OCR-Bench [9]. Finally, because our synthetic dataset was designed to enhance the model’s reasoning capabilities, we chose MMBench [24] and MMMU [46] as additional OOD benchmarks.

### v2

@@ Abstract
Training models on synthetic data has emerged as an increasingly important strategy for improving the performance of generative AI. This approach is particularly helpful for large multimodal models (LMMs) due to the relative scarcity of high-quality paired image-text data compared to language-only data. While a variety of methods have been proposed for generating large multimodal datasets, they do not tailor the synthetic data to address specific deficiencies in the reasoning abilities of LMMs which will be trained with the generated dataset. In contrast, humans often learn in a more efficient manner by seeking out examples related to the types of reasoning where they have failed previously. Inspired by this observation, we propose a new approach for synthetic data generation which is grounded in the analysis of an existing LMM’s reasoning failures. Our methodology leverages frontier models to automatically analyze errors produced by a weaker LMM and propose new examples which can be used to correct the reasoning failure via additional training, which are then further filtered to ensure high quality. We generate a large multimodal instruction tuning dataset containing over 553k examples using our approach and conduct extensive experiments demonstrating its utility for improving the performance of LMMs on multiple downstream tasks. Our results show that models trained on our synthetic data can even exceed the performance of LMMs trained on an equivalent amount of additional real data, demonstrating the high value of generating synthetic data targeted to specific reasoning failure modes in LMMs.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have significantly advanced the state-of-art across a broad range of problem domains, which can be attributed in large part to scaling the size of models and datasets used for training. However, a limiting factor to further scaling of training datasets is the availability of new high-quality data. With recent studies suggesting that we are rapidly approaching saturation of existing public data sources (Villalobos et al. 2022; Xue et al. 2023), synthetic data generation has become an increasingly popular approach for augmenting training datasets. The generation of high-quality synthetic data has become more feasible recently as the capabilities of LLMs have increased, thereby enabling their use for the production of synthetic data at scale (Ding et al. 2024; Yu et al. 2023; Lee et al. 2024; Howard et al. 2022, 2023; Su et al. 2024).
Synthetic data has been especially valuable for training large multimodal models (LMMs), which combine an LLM with a vision encoder to enable text generation conditioned on multimodal inputs. Unlike LLMs, real data for training LMMs is scarce due to the lack of naturally-occurring images paired with high-quality text. Consequently, popular LMMs such as LLaVA (Liu et al. 2023) use partially synthetic data for multimodal instruction tuning, which is often collected by generating synthetic text for real images.
@@
In contrast to the dominant paradigm for generating synthetic training data, humans often learn more efficiently by focusing their attention on examples which are most related to their past reasoning failures. To acquire expert knowledge, Ericsson et al. (1993) argue that humans should “ideally be given explicit instructions about the best method and be supervised by a teacher to allow individualized diagnosis of errors, informative feedback, and remedial part training.” Since humans acquire new knowledge by learning from their failures (Darabi et al. 2018; Brown et al. 1989; Morgan 1904), seeking out additional examples related to past reasoning errors for further practice and evaluation helps the learner achieve expertise in a subject. Problems which require types of reasoning that have already been mastered are often ignored under this paradigm, as they provide little incremental value in the learning process.
@@
Using our approach, we generate a large multimodal instruction tuning dataset containing over 553k examples derived from the reasoning failures of LLaVA-1.5-7B, a popular LMM. We conduct extensive training experiments showing that our synthetic data improves LLaVA’s performance on multiple downstream tasks, even outperforming additional training on alternative real datasets. This is notable, as prior studies have shown that training on synthetic data often requires significantly more examples to achieve equivalent results as training on real data (He et al. 2022). Training on our dataset also produces greater performance gains than previously-proposed synthetic datasets, highlighting the utility of grounding synthetic data generation in the analysis of a model’s reasoning failures.
@@ 2 Related Work
Chen et al. (2024b) proposed the ALLaVA dataset, which consists of 1.3M samples of real images paired with image annotations and question-answer pairs generated by a frontier LMM. While their dataset is similarly suited for multimodal instruction tuning as our dataset, they utilize only real images and do not ground data generation in model failures, whereas our approach can utilize both real and synthetically generated images. Li et al. (2024) generate synthetic question-answer pairs from an LLM for real chart images to address the narrower domain of reasoning-based chart VQA. Yang et al. (2025) generate synthetic multimodal data for text-rich images, leveraging code-guided generation in languages such as LaTeX and HTML to automatically generate synthetic images with dense text descriptions. In contrast, our approach can be broadly applied to different problem domains while leveraging text-to-image diffusion models to produce more diverse types of images.
Other approaches to generating synthetic data suitable for training MLLMs differ from ours in the types of reasoning which they target. Img-Diff (Jiao et al. 2025) contains contrasting synthetic examples which differ only in terms of certain objects in specific regions of an image, requiring the MLLM to learn to identify and articulate such differences. Data-Juicer 2.0 (Chen et al. 2024a) provides multimodal synthetic data generation capabilities utilizing foundation MLLMs and text-to-image diffusion models, but does not address our specific task of generating synthetic data derived from reasoning failures. For a comprehensive data-centric review of MLLMs and how they can assist in the data generation process, see Qin et al. (2025).
@@
Some prior work has explored the use of model failures in the context of synthetic data generation. Jain et al. (2022) proposed an approach to identify directions in a vision model’s latent space which correspond to model failures, which can then be integrated into diffusion models to generate synthetic images tailored to correcting the model’s mistakes. Chegini and Feizi (2023) use ChatGPT and CLIP to identify text descriptions corresponding to a vision model’s failure modes, which are used as prompts for diffusion models to produce synthetic images for data augmentation. In the realm of language-only models, DISCERN (Menon and Srivastava 2024) utilizes two LLMs to iteratively produce natural language descriptions of errors for synthetic data generation. Lee et al. (2024) use incorrect answers from a student LLM finetuned on specific tasks as input to a teacher LLM which generates new examples to use for training. These prior works differ from ours in that they (1) generate only single-modality data (e.g., text-only or image-only), and (2) focus on classification tasks such as image recognition or text classification. In contrast, our approach generates image-text datasets aimed at training models for open-ended text generation conditioned on multimodal inputs.
@@
AgentInstruct (Mitra et al. 2024) is an agentic framework for generating synthetic data from a powerful frontier model (e.g., GPT-4) to teach new skills to a weaker LLM. Similarly, Ziegler et al. (2024) utilize few-shot examples annotated by humans and retrieved documents with produce synthetic data from LLMs for teaching specialized tasks to models. Prompt-based methods for synthetic data generation from LLMs without seed documents (Gupta et al. 2023; Ubani et al. 2023) as well as knowledge distillation from a teacher model (Kim et al. 2024) have also been proposed. Unlike our work, these prior studies focus on language-only data generation and use seed documents (e.g., raw text, source code) or prompts as a basis for data generation rather than an analysis of model failures.
@@ 3.2 Synthetic Data Generation
Using the resulting MFS, we guide the frontier model through a structured multi-turn process to diagnose the failures of the baseline LMM and generate new training samples to address these failure modes.
The frontier model first analyzes the image, question, ground truth, and incorrect prediction to diagnose reasoning errors. It is then instructed to propose new challenging samples consisting of a detailed image description, question, and deterministic answer designed to address the failures. We explore two sourcing strategies for the images: real images and synthetically generated ones. These two methods utilize the prompt described in Figure 5 of the Appendix. For Method 1, the final image-generation step is skipped.
@@
We leverage the original images from failed samples and prompt the frontier model to generate 10 new question-answer pairs per sample using the reasoning pipeline described above. This approach is particularly effective for benchmarks like InfoVQA and ScienceQA, where text-to-image models often fail to capture fine-grained text and generate accurate spatial details.
@@
In this approach, the frontier model generates both a question-answer pair and a detailed image description, which is used to prompt a text-to-image diffusion model to generate synthetic images. For each failed sample, we create 100 fully synthetic samples, consisting of 10 (image prompt, question, answer) triplets, with each image prompt used to produce 10 images using different classifier-free guidance scales to enhance image diversity.
@@
To further enhance the diversity of our data, we use a variation of our prompt which instructs the frontier LMM to “provide examples that challenge Model A’s weaknesses using scenarios from entirely different domains.” This additional instruction is appended to Step 4 (Figure 5) to encourage the generation of samples in different domains and enhance generalization. Figure 2 compares fully synthetic data with similar and non-similar domains. We also varied constraints on the question format (e.g., multiple-choice, true/false) and answer instructions (e.g., requiring ”Unanswerable” or brief responses, see the Shiba Inu example in Figure 10).
@@
Filtering
We use the same frontier LMM to filter synthetic samples by rating image-question-answer triplets from 1 (incorrect) to 3 (fully correct), as shown in the prompt in Figure 6. Only samples rated 3 are retained to ensure quality.
@@ 3.3 Details of Generated Dataset
Our synthetic dataset consists of 553,992 samples incorporating both real and generated images derived from the MFS of LLaVA-1.5-7B on four benchmark training sets: VizWiz (Gurari et al. 2018), InfoVQA (Mathew et al. 2022), ScienceQA (Lu et al. 2022), and OK-VQA (Marino et al. 2019). These benchmarks were chosen to cover a wide range of visual reasoning challenges. VizWiz features real-world images from visually impaired users, demanding detailed scene understanding. OK-VQA requires external knowledge for visual question answering. InfoVQA focuses on text-rich images, testing reading comprehension. ScienceQA involves multimodal scientific questions, assessing spatial and logical reasoning skills. We use GPT-4o (OpenAI 2024) as the frontier LMM for analyzing the reasoning failures of LLaVA-1.5-7B due to its strong multimodal reasoning capabilities. The text-to-image model FLUX.1-schnell (Labs 2024) was used to generate the images with a resolution of 1024×1024 pixels and a classifier guidance scale ranging from 3 to 13.
@@
Table 6 (Appendix) provides an overview of each benchmark, including the size of the training split, the MFS size, and the number of synthetic samples after filtering. For InfoVQA and ScienceQA, we applied Method 1 from Section 3.2. For OK-VQA, we employed Method 2. For the VizWiz, we integrated both methods. The final dataset comprises 42% real-image-based and 58% fully synthetic samples, each with a single-turn question-answer pair.
Our filtering process effectively removes low-quality samples. VizWiz had the highest rejection rate (81% for synthetic images and 34% for real images), highlighting the challenge of generating quality synthetic visuals. OK-VQA’s lower rate (29%) suggests simpler and less ambiguous visual content. ScienceQA showed a similar 29% rejection rate for real images, while InfoVQA had only 5% removed. Figure 3 showcases diverse question types and the quality of generated samples from VizWiz and InfoVQA. Appendix D includes additional examples, including samples with errors or ambiguities. Finally, Figure 8 shows OK-VQA examples with original image-question-answer pairs, the GPT-4o-identified failure mode, and newly generated samples targeting the same failure.
@@ 4.1 Details of Experimental Setting
We perform vision-language instruction tuning with LLaVA-v1.5-7B, building upon the LLaVA-1.5-mix-665K instruction tuning dataset (Liu et al. 2024a). This dataset contains 665K structured user-GPT conversations mostly focused on visual prompts. Because our approach explicitly targets image-grounded reasoning, we only selected the 624K samples that include images from the original LLaVA-instruct data. We then augmented this visually oriented subset with our generated synthetic MFS (Section 3.3), varying the number of synthetic samples used according to the requirements of each experiment. The resulting training set combines real image-text conversations and generated synthetic samples that address specific reasoning failures. For fine-tuning, we used Vicuna-v1.5-7B (Zheng et al. 2023) weights as the LLM backbone and leveraged the pretrained multimodal projector from LLaVA-1.5-7b. We followed the original LLaVA training procedure, ensuring a fair comparison to existing methods (details in Appendix A.5). We refer to models trained with a mixture of the original LLaVA-Instruct dataset and our synthetic data as $\textrm{LLaVA}_{\textrm{syn}}$. As a baseline, we report the performance of LLaVA trained under the same setting but without any additional synthetic data added to the training dataset (i.e., $N_{syn}=0$). Additionally, for each dataset from which reasoning failures were derived for synthetic data generation, we report the performance of a LLaVA model trained on an equivalent amount of real data sourced from the corresponding training dataset (denoted as $\textrm{LLaVA}_{\textrm{real}}$). This provides a measure of the efficiency of our synthetic data relative to training on real in-domain data.